OpenStudy (anonymous):

Calculus question. Just want to check that I did this right. Q. If a snowball melts so that its diameter d decreases at a rate of 2cm/minute, find the rate at which its surface area decreases when the diameter is 20cm. First, I converted the diameter to radius. r=d/2 r=20/2=10 then I took the derivative of the surface area formula. SA=4(pi)(r^2) SA'=4(pi)(r^2)' SA'=8(pi)(r) I then plugged in the radius SA'(at r=10cm)=8(pi)(10) =80(pi) cm/minute Is this right?

OpenStudy (dumbcow):

not quite, you are not factoring in that diameter is decreasing at rate of 2cm/min the rate they want is per time but you took derivative with respect to r not t so you can't just plug in radius use this equation to help \[\frac{dA}{dt} = \frac{dA}{dr}*\frac{dr}{dt}\] by multiplying the 2 rates, the "dr" cancels and you are left with desired rate of surface area per time dr/dt is given as -1 cm/min thus dA/dt = -80pi sq.cm/ min

OpenStudy (anonymous):

thank you :). I got it!